L-1{ s+1 / (s2+2s+s)2 }


1
Expert's answer
2021-09-07T03:15:02-0400
"\dfrac{s+1}{(s^2+2s+s)^2}=\dfrac{s+1}{(s^2+3s)^2}"

"=\dfrac{A}{s}+\dfrac{B}{s^2}+\dfrac{C}{s+3}+\dfrac{D}{(s+3)^2}"

"=\dfrac{As(s+3)^2+B(s+3)^2+Cs^2(s+3)+Ds^2}{(s^2+3s)^2}"

"=\dfrac{As^3+6As^2+9As+Bs^2+6Bs+9B+Cs^3+3Cs^2+Ds^2}{(s^2+3s)^2}"

"s^3:A+C=0"


"s^2:6A+B+3C+D=0"


"s^1:9A+6B=1"


"s^0:9B=1"



"A=\dfrac{1}{27}, B=\dfrac{1}{9}, C=-\dfrac{1}{27}, D=-\dfrac{2}{9}"


"L^{-1}(\dfrac{s+1}{(s^2+2s+s)^2})=\dfrac{1}{27}L^{-1}(\dfrac{1}{s})+\dfrac{1}{9}L^{-1}(\dfrac{1}{s^2})"

"-\dfrac{1}{27}L^{-1}(\dfrac{1}{s+3})-\dfrac{2}{9}L^{-1}(\dfrac{1}{(s+3)^2})"

"=\dfrac{1}{27}+\dfrac{t}{9}-\dfrac{1}{27}e^{-3t}-\dfrac{2}{9}te^{-3t}"



"L^{-1}(\dfrac{s+1}{(s^2+2s+s)^2})=\dfrac{1}{27}+\dfrac{t}{9}-\dfrac{1}{27}e^{-3t}-\dfrac{2}{9}te^{-3t}"